# Zaragoza, Spain Feature Splatting

### v1

@@ Abstract
3D Gaussian Splatting has emerged as a very promising scene representation, achieving state-of-the-art quality in novel view synthesis significantly faster than competing alternatives. However, its use of spherical harmonics to represent scene colors limits the expressivity of 3D Gaussians and, as a consequence, the capability of the representation to generalize as we move away from the training views. In this paper, we propose to encode the color information of 3D Gaussians into per-Gaussian feature vectors, which we denote as Feature Splatting (FeatSplat). To synthesize a novel view, Gaussians are first "splatted" into the image plane, then the corresponding feature vectors are alpha-blended, and finally the blended vector is decoded by a small MLP to render the RGB pixel values. To further inform the model, we concatenate a camera embedding to the blended feature vector, to condition the decoding also on the viewpoint information. Our experiments show that these novel model for encoding the radiance considerably improves novel view synthesis for low overlap views that are distant from the training views. Finally, we also show the capacity and convenience of our feature vector representation, demonstrating its capability not only to generate RGB values for novel views, but also their per-pixel semantic labels. We will release the code upon acceptance.
@@ 1 Introduction
Finding appropriate 3D scene representations is a key challenge in order to enable robotics, VR and AR applications. In general, representations should not only encode a scene’s geometry. Radiance properties are very relevant for novel view synthesis and high-level aspects such as semantics or affordances are essential to define robotics tasks. In the last years, Neural Radiance Fields (NeRFs) [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] have been the dominant approach in the research literature to learn implicit scene representations from images. However, NeRFs are computationally demanding at training and rendering and scale badly with the size of the scene, which has limited their use in practical applications.
@@ 2 Related Work
Recently, 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] has emerged as a substantially faster yet still high-quality alternative. It relies on the explicit modelling of the 3D geometry by a set of 3D Gaussians, each of them with a set of spherical harmonics (SHs) [Yu et al.(2021)Yu, Li, Tancik, Li, Ng, and Kanazawa] to represent their color. Rather than using ray tracing like NeRFs, Gaussians are rasterized by being projected into the image plane, sorted by the Radix algorithm [Merrill and Grimshaw(2010)], and combined using alpha-blending to give the RGB color for every pixel. The rendered image is then used to compute a photometric loss, and the Gaussians’ parameters are optimized using backpropagation. By avoiding the MLP training and relying in an efficient sorting algorithm and CUDA implementation of the rasterizer, 3DGS is the state of the art for the combined rendering performance, training time, and inference rendering speed. Nevertheless, 3DGS is also characterized by 1) a high memory usage; 2) poor surfaces’ reconstructions; and 3) the limited capacity of SHs, which increases the number of Gaussians needed for high quality results, worsening the previous two problems. Several approaches augment Gaussians adding latent vectors which are latter decoded. LangSplat [Qin et al.(2023)Qin, Li, Zhou, Wang, and Pfister] combines CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] features with SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] segmentation masks to generate a language field enabling open-vocabulary 2D semantic segmentation. FMGS [Zuo et al.(2024)Zuo, Samangouei, Zhou, Di, and Li] uses a hash grid based on iNGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] to encode CLIP and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski] features, to also generate an intrinsic representation capable of open-vocabulary 2D segmentation. Nevertheless, none of them neither address the SHs capacity nor show improvement in novel view synthesis, as we do. Based also on iNGP, Compact-3DGS [Lee et al.(2023)Lee, Rho, Sun, Ko, and Park] adds a hash grid that encodes a latent vector for each Gaussian and uses a small MLP conditioned by the viewing direction to recover the corresponding color for each point-of-view. Similarly, our FeatSplat relies on using a latent vector to encode each Gaussian’s color. But, differently from it, we directly attach a feature vector to each Gaussian, and, during the rasterization, rather than first decoding the colors and then alpha-blending, we first alpha-blend the feature vectors of corresponding Gaussians, and then decode the final vector to RGB, which allows us to condition on the intrinsic and extrinsic camera parameters. In our experiments, we show that FeatSplat, due to these differences, outperforms Compact-3DGS at novel view synthesis.
@@ 3 Preliminaries: 3D Gaussian Splatting
Representation. The geometry of a 3D Gaussian $G$ is parameterized by its 3D center position $\mathbf{x}_{G}\in\mathbb{R}^{3}$ and a 3D covariance $\Sigma_{G}=R_{G}S_{G}S_{G}^{\top}R_{G}^{\top}\succeq 0$ that can be decomposed into a rotation matrix $R_{G}\in SO(3)$ and a scaling matrix $S_{G}\in\text{diag}(\mathbb{R}^{3}_{\geq 0})$. Each Gaussian has a density value ${\sigma}_{G}\in\mathbb{R}$; and a set of spherical harmonics (SHs) coeficients $\mathbf{k}_{G}\in\mathbb{R}^{h}$ that parameterize the function $\mathcal{S}_{\mathbf{k}_{G}}\left(\mathbf{d}_{G}\right)$ to encode RGB values $\mathbf{c}_{G}\in\mathbb{R}^{3}$ depending on the viewing direction $\mathbf{d}_{G}\in S^{2}$ relative to each Gaussian.
Differentiable rendering. To render an RGB image $\hat{{I}}\in\mathbb{R}^{w\times h\times 3}$ of the scene, first SHs are converted to RGB, $\mathbf{c}_{G}=\mathcal{S}_{\mathbf{k}_{G}}\left(\mathbf{d}_{G}\right)$. Then, Gaussians are projected, or splatted, into 2D Gaussians in the image plane [Zwicker et al.(2001)Zwicker, Pfister, Van Baar, and Gross] and sorted using the Radix algorithm. Finally, the color $\textbf{c}_{p}$ for each pixel $\mathbf{p}\in\Omega$, $\Omega$ standing for the image domain, comes from the alpha blending of the $\mathcal{N}\subset G$ 2D Gaussians falling into it
@@
Optimization. The representation is fitted as follows. The 3D Gaussians are initialized either randomly or from a sparse Structure from Motion point cloud. At each optimization step, the Gaussians are used to render a viewpoint from the training set and compute the loss
@@ 4 Feature Splatting (FeatSplat)
Semantic segmentation. Learning per-Gaussian feature vectors $\textbf{f}_{p}$ increases the capability of a 3DGS representation. In this paper we illustrate how FeatSplat can be adapted for 2D semantic segmentation. All is needed is changing the output dimension of the MLP from $3$ channels to $3+\#\text{classes}$ and training with an additional semantic cross-entropy loss $\mathcal{L}_{ce}$.
@@ 5 Experiments
Implementation details. We implemented FeatSplat with both $16-$ and $32-$dimensional feature vectors for novel view synthesis, and only $32-$dimensional ones for novel view synthesis plus semantic segmentation. For each scene, we train a MLP with one hidden layer of 64 neurons, common for all Gaussians. For semantic segmentation, the model is trained on 64 classes: the 63 most frequent classes in ScanNet++, plus an unknown category label. Further implementation and optimization details are included in App. A.
Baselines. We compare FeatSplat against 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis], the current state of the art for real-time novel view synthesis, Compact-3DGS [Lee et al.(2023)Lee, Rho, Sun, Ko, and Park] which also augmented 3DGS with learned feature vectors instead of SHs, and iNGP [Müller et al.(2022)Müller, Evans, Schied, and Keller], which was the previous state of the art. Finally, we also include Zip–NeRF and Mip-NeRF 360 metrics on the datasets they have been reported, as a reference. Nevertheless, FeatSplat should not be compared against them due to their offline nature (they both render at less than 1 frame per second).
Datasets and metrics. Rendering performance is evaluated using PSNR, LPIPS [Zhang et al.(2018)Zhang, Isola, Efros, Shechtman, and Wang], and SSIM [Wang et al.(2004)Wang, Bovik, Sheikh, and Simoncelli], computational aspects by the representation size in megabytes (MB), rendering speed in Frames per Second (FPS) on a GTX 3090, and semantic segmentation by the mIOU weighted over classes’ support among all scenes. We evaluate on three common datasets in the literature [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis, Lee et al.(2023)Lee, Rho, Sun, Ko, and Park, Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman]: Mip360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] (T&Ts); and Deep Blending [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow] (DB), for a total of 12 scenes. In these datasets the test set is selected by sampling one out of every eight frames of the image sequences, and the other 7 frames are used to train the model. We also evaluate FeatSplat in a subset of 21 scenes from the novel ScanNet++ [Yeshwanth et al.(2023)Yeshwanth, Liu, Nießner, and Dai]. ScanNet++ is a large–scale dataset with high-resolution RGB and semantic annotations of indoor scenes, and a test set generated from an independent camera trajectory specifically designed to evaluate novel view synthesis with low overlap and separated views.
The independent test set from ScanNet++ represents a considerable increase in difficulty with respect to commonly used datasets. Sampling one out of 8 frames from a video stream (as done in Mip-360, T&Ts and DB) means that at least the previous and following frames would have a significant overlap with the test frames. In contrast, an independent sequence from a different trajectory results in test images looking at unexplored parts of the scene or at the same areas but from rather different viewpoints. In Figure 2 we can visualize, for each of the four datasets, the scene with the highest average angular distance from each test viewpoint (red) to its 10 closest training viewpoints (blue). The contrast between datasets is clear. While on Mip-360, T&Ts, and DB, cameras tend to be nearby each other and looking in the same direction, on ScanNet++ there are several test views that are isolated and looking at mostly unexplored areas. Further details on the datasets can be found in App. B.
@@ 5.1 Novel View Synthesis Results
Out of distribution. In order to analyze points of view far from the training cameras, we sample four points of view from simple linear trajectories on outdoor scenes, see Figure 4. Observe how in Train 3DGS generates floating 3D Gaussians to model far away clouds, which obfuscate the visibility for closer views. Observe instead in FeatSplat–32 that, as the camera gets closer to the train, the Gaussians that before were white, first switch to blue and finally also to green, to better represent the background sky and trees. A similar effect can be observed on the Bicycle scene in the right top corner of the image. Finally, the Treehill data reveals how FeatSplat learned to adapt the Gaussians’ brightness for slightly different positions of the camera while 3DGS did not.
@@
ScanNet++. Table 2 shows that FeatSplat significantly outperforms both 3DGS and Compact-3DGS on all SSIM, PSNR, and LPIPS, with both FeatSplat–16 and FeatSplat–32 achieving the same performance. The gap between FeatSplat and 3DGS is due to a more challenging test set, as previously mentioned. To properly learn view-dependent effects, 3DGS overfits to the training cameras. Some Gaussians are placed in such positions that are seen only in the optimized direction, while not being visible from other training directions without the same effect. This effect is unnoticed from points of view that are close to the training samples, or with a high visual overlap. However, for viewpoints far-away from the training ones, or with lower visual overlap, they become easily noticeable. The ability of FeatSplat to encode multiple colors in the same Gaussian, and to condition the decoding on the point of view, reduces the impact of this problem. Figure 5 shows how FeatSplat renders more accurate colors, captures smaller details, and reduces the number of artifacts.
@@
Semantic FeatSplat Our results so far suggest that 16 dimensions suffice to encode color. We then chose FeatSplat–32 to encode both RGB and semantic labels and perform semantic segmentation. FeatSplat’s segmentation masks for novel views achieve weighted mIOU of $0.629$, maintaining the rendering performance with a PSNR of $24.64$, and SSIM of $0.875$ and an LPIPS of $0.244$, and still rendering in real-time at 56 FPS. The segmentation maps generated, see Figure 6, show that the model produces coherent segmentations even for small elements. However, the edges of the masks are noisy, as they follow Gaussians’ shapes.

### v2

@@ Abstract
3D Gaussian Splatting has emerged as a very promising scene representation, achieving state-of-the-art quality in novel view synthesis significantly faster than competing alternatives. However, its use of spherical harmonics to represent scene colors limits the expressivity of 3D Gaussians and, as a consequence, the capability of the representation to generalize as we move away from the training views. In this paper, we propose to encode the color information of 3D Gaussians into per-Gaussian feature vectors, which we denote as Feature Splatting (FeatSplat). To synthesize a novel view, Gaussians are first "splatted" into the image plane, then the corresponding feature vectors are alpha-blended, and finally the blended vector is decoded by a small MLP to render the RGB pixel values. To further inform the model, we concatenate a camera embedding to the blended feature vector, to condition the decoding also on the viewpoint information. Our experiments show that these novel model for encoding the radiance considerably improves novel view synthesis for low overlap views that are distant from the training views. Finally, we also show the capacity and convenience of our feature vector representation, demonstrating its capability not only to generate RGB values for novel views, but also their per-pixel semantic labels. Code is available on the project page.
@@ 1 Introduction
Finding appropriate 3D scene representations is a key challenge in order to enable robotics, VR and AR applications. In general, representations should not only encode a scene’s geometry. Radiance properties are very relevant for novel view synthesis, and high-level aspects such as semantics or affordances are essential to define robotics tasks. In the last years, Neural Radiance Fields (NeRFs) [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] have been the dominant approach in the research literature to learn implicit scene representations from images. However, NeRFs are computationally demanding at training and rendering and scale badly with the size of the scene, which has limited their use in practical applications.
@@ 2 Related Work
Recently, 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] has emerged as a substantially faster yet still high-quality alternative. It relies on the explicit modelling of the 3D geometry by a set of 3D Gaussians, each of them with a set of spherical harmonics (SHs) [Yu et al.(2021)Yu, Li, Tancik, Li, Ng, and Kanazawa] to represent their color. Rather than using ray tracing like NeRFs, Gaussians are rasterized by being projected into the image plane, sorted by the Radix algorithm [Merrill and Grimshaw(2010)], and combined using alpha-blending to give the RGB color for every pixel. The rendered image is then used to compute a photometric loss, and the Gaussians’ parameters are optimized using backpropagation. By avoiding the MLP training and relying in an efficient sorting algorithm and CUDA implementation of the rasterizer, 3DGS is the state of the art for the combined rendering performance, training time, and inference rendering speed. Nevertheless, 3DGS is also characterized by 1) a high memory usage; 2) poor surfaces’ reconstructions; and 3) the limited capacity of SHs, which increases the number of Gaussians needed for high quality results, worsening the previous two problems. Several approaches augment Gaussians adding latent vectors which are latter decoded. LangSplat [Qin et al.(2023)Qin, Li, Zhou, Wang, and Pfister] combines CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] features with SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] segmentation masks to generate a language field enabling open-vocabulary 2D semantic segmentation. Similarly, both Feature 3DGS [Zhou et al.(2024)Zhou, Chang, Jiang, Fan, Zhu, Xu, Chari, You, Wang, and Kadambi] and Gaussian Grouping [Ye et al.(2024)Ye, Danelljan, Yu, and Ke], use SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] as a teacher to embed in each Gaussian a semantic feature. FMGS [Zuo et al.(2024)Zuo, Samangouei, Zhou, Di, and Li] uses a hash grid based on iNGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] to encode CLIP and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski] features, to also generate an intrinsic representation capable of open-vocabulary 2D segmentation. Nevertheless, none of them neither address the SHs capacity nor show improvement in novel view synthesis, as we do. Based also on iNGP, Compact-3DGS [Lee et al.(2023)Lee, Rho, Sun, Ko, and Park] adds a hash grid that encodes a latent vector for each Gaussian and uses a small MLP conditioned by the viewing direction to recover the corresponding color for each point-of-view. Similarly, our FeatSplat relies on using a latent vector to encode each Gaussian’s color. But, differently from it, we directly attach a feature vector to each Gaussian, and, during the rasterization, rather than first decoding the colors and then alpha-blending, we first alpha-blend the feature vectors of corresponding Gaussians, and then decode the final vector to RGB, which allows us to condition on the intrinsic and extrinsic camera parameters. In our experiments, we show that FeatSplat, due to these differences, outperforms Compact-3DGS at novel view synthesis.
@@ 3 Preliminaries: 3D Gaussian Splatting
Representation.
The geometry of a 3D Gaussian $G$ is parameterized by its 3D center position $\mathbf{x}_{G}\in\mathbb{R}^{3}$ and a 3D covariance $\Sigma_{G}=R_{G}S_{G}S_{G}^{\top}R_{G}^{\top}\succeq 0$ that can be decomposed into a rotation matrix $R_{G}\in SO(3)$ and a scaling matrix $S_{G}\in\text{diag}(\mathbb{R}^{3}_{\geq 0})$. Each Gaussian has a density value ${\sigma}_{G}\in\mathbb{R}$; and a set of spherical harmonics (SHs) coeficients $\mathbf{k}_{G}\in\mathbb{R}^{h}$ that parameterize the function $\mathcal{S}_{\mathbf{k}_{G}}\left(\mathbf{d}_{G}\right)$ to encode RGB values $\mathbf{c}_{G}\in\mathbb{R}^{3}$ depending on the viewing direction $\mathbf{d}_{G}\in S^{2}$ relative to each Gaussian.
Differentiable rendering.
To render a RGB image $\hat{{I}}\in\mathbb{R}^{w\times h\times 3}$ of the scene, first SHs are converted to RGB, $\mathbf{c}_{G}=\mathcal{S}_{\mathbf{k}_{G}}\left(\mathbf{d}_{G}\right)$. Then, Gaussians are projected, or splatted, into 2D Gaussians in the image plane [Zwicker et al.(2001)Zwicker, Pfister, Van Baar, and Gross] and sorted using the Radix algorithm. Finally, the color $\textbf{c}_{p}$ for each pixel $\mathbf{p}\in\Omega$, $\Omega$ standing for the image domain, comes from the alpha blending of the $\mathcal{N}\subset G$ 2D Gaussians falling into it
@@
Optimization.
The representation is fitted as follows. The 3D Gaussians are initialized either randomly or from a sparse Structure from Motion point cloud. At each optimization step, the Gaussians are used to render a viewpoint from the training set and compute the loss
@@ 4 Feature Splatting (FeatSplat)
Semantic segmentation.
Learning per-Gaussian feature vectors $\textbf{f}_{p}$ increases the capability of a 3DGS representation. In this paper we illustrate how FeatSplat can be adapted for 2D semantic segmentation. All is needed is changing the output dimension of the MLP from $3$ channels to $3+\#\text{classes}$ and training with an additional semantic cross-entropy loss $\mathcal{L}_{ce}$.
@@ 5 Experiments
Implementation details.
We implemented FeatSplat with both $16-$ and $32-$dimensional feature vectors for novel view synthesis, and only $32-$dimensional ones for novel view synthesis plus semantic segmentation. For each scene, we train a MLP with one hidden layer of 64 neurons, common for all Gaussians. For semantic segmentation, the model is trained on 64 classes: the 63 most frequent classes in ScanNet++, plus an unknown category label. Further implementation and optimization details are included in App. A.
Baselines.
We compare FeatSplat against 3DGS [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis], the current state of the art for real-time novel view synthesis, Compact-3DGS [Lee et al.(2023)Lee, Rho, Sun, Ko, and Park] which also augmented 3DGS with learned feature vectors instead of SHs, and iNGP [Müller et al.(2022)Müller, Evans, Schied, and Keller], which was the previous state of the art. Finally, we also include Zip–NeRF and Mip-NeRF 360 metrics on the datasets they have been reported, as a reference. Nevertheless, FeatSplat should not be compared against them due to their offline nature (they both render at less than 1 frame per second). Finally, on the semantic segmentation experiment, we do not compare against other models that learn specific semantic features for open-vocabulary semantic segmentation. Given that our model is trained on a closed-vocabulary, we would have an unfair advantage.
@@
Datasets and metrics.
Rendering performance is evaluated using PSNR, LPIPS [Zhang et al.(2018)Zhang, Isola, Efros, Shechtman, and Wang], and SSIM [Wang et al.(2004)Wang, Bovik, Sheikh, and Simoncelli], computational aspects by the representation size in megabytes (MB), rendering speed in Frames per Second (FPS) on a GTX 3090, and semantic segmentation by the mIOU weighted over classes’ support among all scenes. We evaluate on three common datasets in the literature [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis, Lee et al.(2023)Lee, Rho, Sun, Ko, and Park, Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman]: Mip360 [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman], Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] (T&Ts); and Deep Blending [Hedman et al.(2018)Hedman, Philip, Price, Frahm, Drettakis, and Brostow] (DB), for a total of 12 scenes. In these datasets the test set is selected by sampling one out of every eight frames of the image sequences, and the other 7 frames are used to train the model. We also evaluate FeatSplat in a subset of 21 scenes from the novel ScanNet++ [Yeshwanth et al.(2023)Yeshwanth, Liu, Nießner, and Dai]. ScanNet++ is a large–scale dataset with high-resolution RGB and semantic annotations of indoor scenes, and a test set generated from an independent camera trajectory specifically designed to evaluate novel view synthesis with low overlap and separated views.
The independent test set from ScanNet++ represents a considerable increase in difficulty with respect to commonly used datasets. Sampling one out of 8 frames from a video stream (as done in Mip-360, T&Ts and DB) means that at least the previous and following frames would have a significant overlap with the test frames. In contrast, an independent sequence from a different trajectory results in test images looking at unexplored parts of the scene or at the same areas but from rather different viewpoints. In Figure 2 we can visualize, for each of the four datasets, the scene with the highest average angular distance from each test viewpoint (red) to its 10 closest training viewpoints (blue). The contrast between datasets is clear. While on Mip-360, T&Ts, and DB, cameras tend to be nearby each other and looking in the same direction, on ScanNet++ there are several test views that are isolated and looking at mostly unexplored areas. Further details on the datasets can be found in App. C.
@@ 5.1 Novel View Synthesis Results
Figure 5: Global illumination change modifying MLP input values of the pixel embedding (PE), and the 3D coordinates (X, Y, Z), on scene 03f7a0e617 from ScanNet++.
@@
Out of distribution.
In order to analyze points of view far from the training cameras, we sample four points of view from simple linear trajectories on outdoor scenes, see Figure 4. Observe how, in Train, 3DGS generates floating 3D Gaussians to model far away clouds, which obfuscate the visibility for closer views. Observe instead in FeatSplat–32 that, as the camera gets closer to the train, the Gaussians that before were white, first switch to blue and finally also to green, to better represent the background sky and trees. A similar effect can be observed on the Bicycle scene in the right top corner of the image. Finally, the Treehill data reveals how FeatSplat learned to adapt the Gaussians’ brightness for slightly different positions of the camera while 3DGS did not.
Light encoding.
Furthermore, we analyze the pixel and position embeddings impact on the MLP by individually evaluating different input values for each of them, see Figure 5. Apparently, the MLP learns to use these embeddings to codify global lightning information. As a consequence, our model allows to modify the scene lightning at inference time without any addition training.
@@
ScanNet++.
Table 2 shows that FeatSplat significantly outperforms both 3DGS and Compact-3DGS on all SSIM, PSNR, and LPIPS, with both FeatSplat–16 and FeatSplat–32 achieving the same performance. The gap between FeatSplat and 3DGS is due to a more challenging test set, as previously mentioned. To properly learn view-dependent effects, 3DGS overfits to the training cameras. Some Gaussians are placed in such positions that are seen only in the optimized direction, while not being visible from other training directions without the same effect. This effect is unnoticed from points of view that are close to the training samples, or with a high visual overlap. However, for viewpoints far-away from the training ones, or with lower visual overlap, they become easily noticeable. The ability of FeatSplat to encode multiple colors in the same Gaussian, and to condition the decoding on the point of view, reduces the impact of this problem. Figure 6 shows how FeatSplat renders more accurate colors, captures smaller details, and reduces the number of artifacts.
Semantic FeatSplat
Our results so far suggest that 16 dimensions suffice to encode color. We then chose FeatSplat–32 to encode both RGB and semantic labels and perform semantic segmentation. FeatSplat’s segmentation masks for novel views achieve weighted mIOU of $0.629$, maintaining the rendering performance with a PSNR of $24.64$, and SSIM of $0.875$ and an LPIPS of $0.244$, and still rendering in real-time at 56 FPS. The segmentation maps generated, see Figure 7, show that the model produces coherent segmentations even for small elements. However, the edges of the masks are noisy, as they follow Gaussians’ shapes. Nevertheless, it proves the versatility of these features and their potential to encode different kind of information.